In a parabola show that the tangent at any point makes equal angles with the foci radius of the point and the line parallel to the x-axis through the point
1
Expert's answer
2020-06-17T18:51:28-0400

We begin with a parabola and choose the coordinate system so that the equation takes the form "x^2=4cy, c>0." We can express "y" in terms of "x" by writing


"y={x^2\over 4c}"

Since "\dfrac{dy}{dx}=\dfrac{x}{2c}," the tangent line at the point "P(x_0,y_0)" has slope "m=\dfrac{x_0}{2c}" and has equation 


"y-y_0=\dfrac{x_0}{2c}(x-x_0)"

The focus is "F(0, c)."

Since the tangent line at "P(x_0, y_0)" is not vertical, it intersects the y-axis at some point "T." To find the coordinates of T, we set "x=0" in the equation of the tangent line


"y-y_0=\dfrac{x_0}{2c}(0-x_0)"

"y=y_0-\dfrac{x_0^2}{2c},\ y_0=\dfrac{x_0^2}{4c},"

"y=-y_0=-\dfrac{x_0^2}{4c}"

We have point "T(0,-y_0)" and "d(F, T)=c+y_0."



Find the distance "d(F, P)"


"d(F,P)=\sqrt{(x_0-0)^2+(y_0-c)^2}="

"=\sqrt{4cy_0+y_0^2-2y_0c+c^2}=\sqrt{(y_0+c)^2=}"

"=|y_0+c|=y_0+c=d(F,T)\ (c>0, y_0>)"

Since "d(F,P)=d(F,T)," the triangle "TFP" is isosceles and the angles marked "\alpha" and "\beta" are equal. Since a ray "l" is parallel to the y-axis, "\alpha=\gamma" and thus "\beta=\gamma)."

This means that light emitted from a source at the focus of a parabolic mirror is reflected in a beam parallel to the axis of that mirror; this is the principle of the searchlight.